OpenStudy (anonymous):

find the differential equation for the DE y=alpha(e^x) +beta(e^2X)+delta(e^3x)

OpenStudy (unklerhaukus):

\[y=\alpha e^x+\beta e^{2x}+\delta e^{3x}\]

OpenStudy (anonymous):

exactly ...jst that I could not write it in that way on the questions!!

OpenStudy (anonymous):

let me give u the answer maybe u will be able to help out!!

OpenStudy (unklerhaukus):

are alpha, beta, delta functions of x?

OpenStudy (anonymous):

y'''-6y''+11y'-6y=0 this is the answer given on the memo

OpenStudy (unklerhaukus):

what was the exact question

OpenStudy (anonymous):

Find a differential equation for the given family of curves... that was one of the equations!!

OpenStudy (unklerhaukus):

so there are three constants , we must take the derivative three times \[y=\alpha e^x+\beta e^{2x}+\delta e^{3x}\] \[y'=\alpha e^x+2\beta e^{2x}+3\delta e^{3x}\]\[y''=\alpha e^x+4\beta e^{2x}+9\delta e^{3x}\]\[y'''=\alpha e^x+8\beta e^{2x}+27\delta e^{3x}\]

OpenStudy (unklerhaukus):

i dont know

OpenStudy (anonymous):

nd then... ow cool..let's jst wait maybe se1 else might crack it!

OpenStudy (unklerhaukus):

@mukushla

OpenStudy (anonymous):

characteristic eqn. of the DE has the eigenvalues in the solution. You can expand (lambda -1)(lambda-2)(lamdba-3) to find the coefficients of the DE.

OpenStudy (anonymous):

am still confused.... coz I have not done eigenvalues yet!!

OpenStudy (experimentx):

\[ (D-1)(D-2)(D-3) y = 0\] the standard solution of homogeneous constant coefficient DE

OpenStudy (experimentx):

besides that ... you can eliminate the value of alpha, beta, delta from those three linear systems ... and put it into the original ... another system and get those values.

OpenStudy (anonymous):

so I eliminate by using the 3 equations of derivatives?that sounds relevent?

OpenStudy (experimentx):

yeah ... that is one of not so nice way that will always get you answer.

OpenStudy (anonymous):

so the experiment thanks dawg I use the standard solution method and it's was just 2minute nd done!!